The shelf life of a particular dairy product is normally distributed with a mean of 12 days and a standard deviation of 3 days. What percent of the product that between 10 and 15 days?

1
Expert's answer
2022-03-18T00:54:36-0400

"P(10<X<15)=P(\frac{10-12}{3}<Z<\frac{15-12}{3})=P(-0.67<Z<1)="

"=P(Z<1)-P(Z<-0.67)=0.5899=58.99\%."